OpenStudy (anonymous):

help me simplify 120^3/sqrt 16^9

OpenStudy (shamim):

\[\frac{ 120 }{ \sqrt{16^{9}} }=\frac{ 120 }{ 4^{9} }\]

OpenStudy (anonymous):

thnx

OpenStudy (shamim):

welcome

OpenStudy (shamim):

anyway may i know ur country name

OpenStudy (anonymous):

ummmmm wat?

OpenStudy (anonymous):

u.s.a

OpenStudy (shamim):

ok